The sum of the page numbers on consecutive pages of a book is 267.what are the page numbers? I had tried this problem but i wasn't understanding it at all.so this is what i tried to do:x+(x+1)=267,2x+1=267,2x=266.My answers were 133, and 134.
Answer by checkley77(12569)

I get the same page numbers as follows:
x+x+1=267
2x=267-1
2x=266
x=266/2
x=133 ans. for the left hand page.
133+1=134 ans. for the right hand page.